Solve by substitution:
4x+y=10
14x-y=-1

Answer by mananth(16946)

4 x + 1 y = 10 .............1
1 y = 10 + -4 x
/ 1 =
y = 10 + -4 x
14 x + -1 y = -1 .....................2
Substitute y in (2)
14 x + -1 ( 10 + -4 x ) =
14 x + -10 + 4 x = -1
14 x 4 x = 9
18 x = 9
/ 18
x= 1/ 2
Plug the value of x in (1)
4 x + 1 y = 10
4 * 1/ 2 + 1 y = 10
2 + 1 y = 10
1 y = 8
/ 1
y= 8
m.ananth@hotmail.ca